# Tweezer interferometry with NOON states

### v1

@@ IX Discussion
An optional way to perform phase estimation is to use the non-linear splitter in reverse. In such a procedure, the states |+ket|$\rangle$| + ⟩ and |ket|$\rangle$| - ⟩ adiabatically become $|N,0\rangle$ and $|N{-}1,1\rangle$. The difference between the outcome states is microscopic, and hence difficult to detect. The natural suggestion would be to use two-level-dynamics that maps |+ket|$\rangle$| + ⟩ and |ket|$\rangle$| - ⟩ to $|N,0\rangle$ and $|0,N\rangle$ respectively. The latter are easily distinguished. The problem with this method is its slowness. Exponentially long time is required because the coupling between the two states is exponentially small in $N$, namely, ${J_{\text{eff}}=Je^{-cN}}$.
Our suggested procedure for phase-estimation is to reverse the splitting, and then to sweep down $J$ but with a non-zero bias. To execute this process, one needs a good experimental control over $\Delta$ to allow ramping it to a small final non-zero value. This maps the $|N,0\rangle$ and $|N{-}1,1\rangle$ states into the macroscopically distinct states $|N,0\rangle$ and $|0,N\rangle$ respectively. Importantly, the required time to reach the desired outcome scales as $N$ and not as $e^{cN}$.
From the implementation perspective, a pertinent question emerges: how can we distinguish between splitting, branching, and spontaneous symmetry breaking? For example, in the experiment of [25], spontaneous symmetry breaking was reported and studied. As we have demonstrated, a minor imbalance $\Delta$ between the two wells can indeed lead to spontaneous symmetry breaking. However, we have found that the required level of balance between the wells scales favorably with the number of particles. Our feasibility analysis has shown that with approximately 50 atoms, the imbalance should be kept below $10^{-6}$, which is challenging but achievable. Even if such a level is attained in the experiment, the only way to verify coherent splitting is to perform the full interferometry scheme and observe $\varphi$-dependent output.
@@ Appendix A The characteristic frequencies
where $\Theta$ is the step function. Here we use the conjugate coordinates $(n,\phi)$ relative to the East pole, meaning that ${n=0}$ indicates ${S_{x}=N/2}$. We use units such that formally ${\hbar=1}$, but in the WKB context this statement needs some clarification. The WKB quantization condition is
@@
where $h$ is the area of Planck cell. For the Bloch sphere it is implicit that the total area of phase space is $2\pi N=(N+1)h$. Form this follows that ${h=2\pi N/(N+1)\approx 2\pi}$. Note that the WKB quantization condition remains the same if we look on the complementary area, taking the West pole as the origin. The characteristic frequency is calculated as follows:
@@
We take $E_{x}=-JN/2$ as the reference for the energy. This is the minimum energy for $J>J_{c}$ or the seperatrix energy for $J<J_{c}$. Accordingly we substitute ${E\mapsto E_{x}+E}$, assume for simplicity ${E>E_{x}}$, and get the following integral expression
$A^{\prime}(E)=\ \ \frac{d}{dE}\left[\int_{-\pi}^{\pi}n_{+}(E)d\phi\right]$
$\ \ \ \ =4\int_{0}^{\pi/2}\!\!\!\frac{d\phi}{\sqrt{(J{-}NU\sin^{2}\phi)^{2}+4EU\sin^{2}\phi}}$ (33)
@@
Adding the the contribution together, the arbitrary borders $b$ and $c$ cancel out, and we get

### v2

@@ VIII Discussion
An optional way to perform phase estimation is to use the non-linear splitter in reverse. In such a procedure, the states |+ket|$\rangle$| + ⟩ and |ket|$\rangle$| - ⟩ adiabatically become the original ground-state condensation and its single-particle excitation. The difference between the outcome states is microscopic, and hence difficult to detect. The natural suggestion would be to use two-level-dynamics that maps |+ket|$\rangle$| + ⟩ and |ket|$\rangle$| - ⟩ to $|N,0\rangle$ and $|0,N\rangle$ respectively. The latter are easily distinguished. The problem with this method is its slowness. Exponentially long time is required because the coupling between the two states is exponentially small in $N$, namely, ${J_{\text{eff}}=Je^{-cN}}$. We note that there are also more exotic suggestions as in [38] that require to go beyond the two-orbital picture.
Our preferred procedure for phase-estimation is to reverse the splitting, and then to sweep down $J$ but with a non-zero bias. To execute this process, one needs a good experimental control over $\Delta$ to allow ramping it to a small final non-zero value. This maps the $\nu=1$ and $\nu=2$ states into the macroscopically distinct states $|N,0\rangle$ and $|0,N\rangle$ respectively. Importantly, the required time to reach the desired outcome scales as $N$ and not as $e^{cN}$.
From the implementation perspective, a pertinent question emerges: how can we distinguish between splitting, branching, and spontaneous symmetry breaking? For example, in the experiment of [39], spontaneous symmetry breaking [40] was reported and studied. As we have demonstrated, a minor uncontrolled imbalance $\Delta$ between the two wells during the splitting phase can indeed lead to spontaneous symmetry breaking. However, we have found that the required level of balance between the wells scales favorably with the number of particles. Our feasibility analysis has shown that with approximately 50 atoms, the imbalance should be kept below $10^{-6}$, which is challenging but achievable. Even when such a level is attained in the experiment, the only way to verify unambiguously coherent splitting is to perform the full interferometry scheme and observe $\varphi$-dependent output.
@@ Appendix A The characteristic frequencies
where $\Theta$ is the step function. Here we use the conjugate coordinates $(n,\phi)$ relative to the X direction, meaning that ${n=0}$ indicates ${S_{x}=N/2}$. We use units such that formally ${\hbar=1}$, but in the WKB context this statement needs some clarification. The WKB quantization condition is
@@
where $h$ is the area of Planck cell. For the Bloch sphere it is implicit that the total area of phase space is $2\pi N=(N+1)h$. Form this follows that ${h=2\pi N/(N+1)\approx 2\pi}$. Note that the WKB quantization condition remains the same if we look on the complementary area, taking ${S_{x}=-N/2}$ as the origin. The characteristic frequency is calculated as follows:
@@
We take $E_{x}=-JN/2$ as the reference for the energy. This is the minimum energy for $J>J_{c}$ or the separatrix energy for $J<J_{c}$. Accordingly we substitute ${E\mapsto E_{x}+E}$, assume for simplicity ${E>E_{x}}$, and get the following integral expression
$A^{\prime}(E)=\ \ \frac{d}{dE}\left[\int_{-\pi}^{\pi}n_{+}(E)d\phi\right]$ (35)
$\ \ \ \ =4\int_{0}^{\pi/2}\!\!\!\frac{d\phi}{\sqrt{(J{-}NU\sin^{2}\phi)^{2}+4EU\sin^{2}\phi}}$
@@
Adding the contributions together, the arbitrary borders $b$ and $c$ cancel out, and we get
@@
Appendix B Crossover to diabatic scenario
[…]
In the adiabatic range, the leakage $\left\langle n_{\text{ex}}\right\rangle$ to higher levels is zero, because the final state is a superposition of the $S_{z}=-N/2$ and $S_{z}=N/2$ states for which $n_{\text{ex}}=0$. In the diabatic range, the sudden approximation implies that the state stays at $S_{x}=N/2$, therefore $n_{\text{ex}}\sim N/2$.